General Physics

the resultant of two forces is root under 41n when they act at right angle and root under 61 n when they act at an angle of 60 degree . find the forces

rahul
11 Years agoGrade

1 Answer

Ravi
11 Years ago
The resultant of 2 forces at an angle theta is
[A2+B2+2AB Cos(Theta)]½

Using this,
When angle is 90
A2+B2= 41 ...(i)

And when angle is 60,
A2+B2+2AB cos60=61 ...(ii)
A2+B2+2AB(1/2)=61
From i
AB=20 ...(iii)
So, Put iii in ii to get the value of A and then B